OpenStudy (anonymous):

1.Angelo took a group of people to the movies and spent $80 on tickets. Adult tickets cost $10, and child tickets cost $6. The equation to represent his cost is 10A + 6C = $80. If there were 5 children in the group, how many adult tickets did he purchase? Show your work and use a complete sentence to explain your solution in the context of the original problem. 2.A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving t

OpenStudy (anonymous):

1. Since you have the equation 10A+6C=80. Then the A represents the number of adults and the C represents the number of children. So to see how many adult tickets he purchased you plug 5 into the c and get: 80=10A+6*5 80=10A+30 Then solve for A 80-30=10A 50=10A A=5. So 5 adult tickets were bought 2. To solve equations with absolute value you have to have think about both the positive and negative and solve for both so: (x-.525)=.01 and -(x-.525)=.01 To solve (x-.525)=.01 add .525 to both sides to get x=.535 To solve -(x-.525)=.01 first distribute the negative -x+.525=.01 subtract .525 from both sides -x=-.515 then divide both sides by -1 x=.515 The answer is x=.535 or x=.515